Problem 9-3A a-b Sheridan Limited purchased a machine on account on April 2, 2018, at an invoice price of $369,950. On AprilCalculate the annual depreciation and total depreciation over the assets life using: (Round the depreciation cost per unit t(2) Double-diminishing-balance method Year Opening Carrying Amount Depreciation Expense Accumulated Depreciation Carrying Amo(3) Units-of-production method Year Units-of-production Depreciation Expense Accumulated Depreciation Carrying Amount 2018 $
Answer #1
Computation of Cost of Machinery
Invoice Price $369,950.00
Delivery Expense $1,900.00
Installation and testing $7,790.00
Cost of Machienry $379,640.00
Straight-Line Depreciation Schedule
Depreciation for the Year
Asset Depreciable Useful Life Depreciation Accumulated Book
year Cost Cost Expense Depreciation Value
2018 $379,640.00
2018 $301,170.00 5 $40,156.00 $40,156.00 $339,484.00
(301170/5*8/12)
2019 $301,170.00 /                     5 = $60,234.00 $100,390.00 $279,250.00
2020 $301,170.00 /                     5 = $60,234.00 $160,624.00 $219,016.00
2021 $301,170.00 /                     5 = $60,234.00 $220,858.00 $158,782.00
2022 $301,170.00 /                     5 = $60,234.00 $281,092.00 $98,548.00
2023 $301,170.00 /                     5 = $20,078.00 $301,170.00 $78,470.00
(301170/5*4/12)
Depreciable cost =379640-78470=$301170
Depreciation per Unit = (Original Cost- Salvage Value)/ Estimated Unit
=(379640-78470)/6141 Unit=$49.0425
Units-of-Production Depreciation Schedule
Depreciation for the Year
Asset Depreciation Number of Depreciation Accumulated Book
Date Cost Per Unit Unit Expense Depreciation Value
2018 $379,640
2018 49.0425 x 900 = $44,138 $44,138 $335,501.75
2019 49.0425 x 1537 = $75,378 $119,517 $260,123.43
2020 49.0425 x 1495 = $73,319 $192,835 $186,804.89
2021 49.0425 x 1174 = $57,576 $250,411 $129,229.00
2022 49.0425 x 1035 = $50,759 $301,170 $78,470.01
Double-Declining-Balance Depreciation Schedule
Depreciation for the Year
Asset Book DDB (1/5) Depreciation Accumulated Book
Date Cost Value Rate Expense Depreciation Value
2018 $379,640.00
2018 $379,640.00 X 20% $50,618.67 $50,618.67 $329,021.33
(379640*20%*8/12)
2019 $329,021.33 X 20% $65,804.27 $116,422.93 $263,217.07
2020 $263,217.07 X 20% $52,643.41 $169,066.35 $210,573.65
2021 $210,573.65 X 20% $42,114.73 $211,181.08 $168,458.92
2022 $168,458.92 X 20% $33,691.78 $244,872.86 $134,767.14
2023 $134,767.14 X 20% $56,297 $301,170.00 $78,470.00
(134767.14-78470)